Calculate the framework electrons using Wade's rule B6H6-2

1
Expert's answer
2021-01-07T05:31:30-0500

Total number valence electrons;

"=(5\times B)+(6\times H)+2"

"=(6\times 3)+(6\times 1)+(2)"

"=18+6+2=26"

Number of electrons for each "B-H" unit"=(6\times 2)=12"

Number of skeletal electrons "=26-12=14"

Number of "SEP=" "14\over 2" "=7"

If "n+1=7" and boron atoms, then "n=6"